Data Reification

(AccId ↪ (2^{AccHolder} × Amount))

⊴_{33} ((AccId ↪ 2^{AccHolder}) × (AccId ↪ Amount))

> abs = join

> rep = nJoinInv

⊴_{24} ((2^{AccHolder} × 1)^{AccId} × (AccId ↪ Amount))

> abs = prod(?, id)

> rep = prod(?, id)

⊴_{1} ((AccId ↪ Amount) × (2^{AccHolder} × 1)^{AccId})

> abs = const(p2, p1)

> rep = const(p2, p1)

⊴_{38} (2^{(AccId×Amount)} × (2^{AccHolder} × 1)^{AccId})

> abs = prod(mkf, id)

> rep = prod(mkr, id)

⊴_{25} (2^{(AccId×Amount)} × (AccId ↪ 2^{AccHolder}))

> abs = prod(id,?)

> rep = prod(id,?)